Example 3: Draw a flowchart that performs the following: Ask a user to enter a number. If the number is between 0 and 10, write the word blue. If the number is between 10 and 20, write the word red. if the number is between 20 and 30, write the word green. If it is any other number, write that it is not a correct color option.

Example 4: Draw a flowchart to print all multiples of 5 between 1 and 100 (including both 1 and 100).

Example 5: Draw a flowchart that will count all the even numbers up to a user defined stopping point.

Example 6: Draw a flowchart that will perform the following.

a) Read in 5 separate numbers.

b) Calculate the average of the five numbers.

c) Find the smallest (minimum) and largest (maximum) of the five entered numbers.

d) Write out the results found from steps b and c with a message describing what they are
Answer #1

Question 3:

Start Enter Number No No Yes If (O<Number <10) If (20<Number<30) Print(Red) Yes Print (Green Print(not a correct option) Stop

Question 4:

Start Input Num No If Num mod 5 =0 Yes Print Num Stop

Question 5:

Start Read Count Set X to 0 Stop No X < Count Print Even Yes X=X+1 Set even = even+2

As per chegg guidelines for multiple questions we will provide answer for first question mandatory. I provide answer for first 3 questions. So to get answer for other questions you'll be able to post it again.